OpenStudy (anonymous):

POLYNOMIALS: SOLVING FOR ROOTS AND UNKOWN COEFFICIENT ....help please! (equation is posted below)

OpenStudy (anonymous):

"Two roots of: \( \Huge\color{blue} {2x^3 +13x^2+22x+k=0} \) are reciprocals of eachother. Find k and all roots of the equation.

ganeshie8 (ganeshie8):

latex addict :P

OpenStudy (anonymous):

Shhh...

OpenStudy (anonymous):

hahaha

ganeshie8 (ganeshie8):

say, the roots are : \(\alpha, \dfrac{1}{\alpha}, \beta\)

OpenStudy (anonymous):

ok then \[P()=(x-\alpha) (x-\beta)(x \alpha-1)\]

ganeshie8 (ganeshie8):

\(\large 2x^3 + 13x^2 + 22x+k = 2(x-\alpha)(x-\dfrac{1}{\alpha})(x-\beta)\)

OpenStudy (anonymous):

Ooohh... ok so I have three unknows therefore i need three equations?

ganeshie8 (ganeshie8):

comparing constant terms both sides, you wud get : \(\large -2\alpha \dfrac{1}{\alpha} \beta = k\)

OpenStudy (anonymous):

so -2b=K ...1

ganeshie8 (ganeshie8):

that gives you : \(\large k = -2 \beta\)

OpenStudy (anonymous):

yep gotcha :)

OpenStudy (anonymous):

Could I do P(b) =0 to find b or ...?

ganeshie8 (ganeshie8):

compare x coefficients both sides : \(\large 22= -2(\alpha + \dfrac{1}{\alpha} + \beta ) \) compare x^2 coefficients both sides : \(\large 13 = 2(1 + \dfrac{\beta}{\alpha} + \alpha \beta)\)

ganeshie8 (ganeshie8):

you can solve these easily for \(\alpha\) and \(\beta\)

OpenStudy (anonymous):

ooohh oK, im trying it now, THANKS @ganeshie8 !!

ganeshie8 (ganeshie8):

maybe il get u started a bit, wait it can be tricky if u dont start correctly..

OpenStudy (anonymous):

hmm, for your second equation I got \[13=2(1+\frac{ \beta }{ \alpha }- \alpha \beta)\] not PLUS ab...

OpenStudy (whpalmer4):

Hmm, I'm not sure those equations are correct. I've solved the problem and my values of alpha and beta satisfy the constraints but not those equations...

ganeshie8 (ganeshie8):

You have below equations to solve : \(\large 13= -2(\alpha + \dfrac{1}{\alpha} + \beta ) \) ---------(1) \(\large 22= 2(1 + \dfrac{\beta}{\alpha} + \alpha \beta) \) -------(2) from the equaiton (1), \(\large \alpha + \dfrac{1}{\alpha} = -\dfrac{13}{2} - \beta\) plug this value in equaiton (2) : \(\large 22= 2(1 + \beta(-\dfrac{13}{2}-\beta))\) you can sovle this, just a quadratic

ganeshie8 (ganeshie8):

ahh flipped x^2 and x coefficients earlier, corrected now.. thanks @whpalmer4 :) plz check if they're ok now :)

OpenStudy (anonymous):

@ganeshie8 I got negatives every time you got a positive for the above equations... @whpalmer4 can you see if your values satisfy the equations when there is a negative""?

OpenStudy (whpalmer4):

note that there are two solutions to that quadratic, and one of them gives you a number that doesn't give you reciprocal roots :-)

OpenStudy (anonymous):

which quadratic are you referring to @whpalmer4 ?

OpenStudy (whpalmer4):

\[\large 22= 2(1 + \beta(-\dfrac{13}{2}-\beta))\]

OpenStudy (whpalmer4):

the "wrong" value corresponds to a value of \(k=5\)

OpenStudy (whpalmer4):

the "right" value gives you roots of the big polynomial which are all multiples of each other (and so for that matter is \(k\))

OpenStudy (anonymous):

Ok, I'll have a go myself and I'll let you know if I have any troube- thanks for your help @ganeshie8 and @whpalmer4 !!! I really appreciate your time :)

OpenStudy (anonymous):

Since \(k/2\) is the product of three roots and \(1\) is the product of two of them (they are reciprocals), we can claim that \(k/2\) is a root.

OpenStudy (whpalmer4):

A different approach, if you're going to use Wolfram technology is just to ask to have the various possible polynomials factored :-) it turns out few negative values of \(k\) give you a factorable polynomial, but there are a number of positive values which will give you a polynomial with rational roots. 0,3,5,8,11,18,35, 96, 143, etc. Yes, \(k/2\) is indeed a root, and that's a good observation.

OpenStudy (whpalmer4):

actually, it's not a perfect observation, though, because the sign is wrong :-)

ganeshie8 (ganeshie8):

hahah im scratching my head around whether complex root pair can be reciprocals of each other

OpenStudy (anonymous):

Can anyone explain why \(\dfrac{k}{2}\) is not the product of roots?

ganeshie8 (ganeshie8):

when k = 18, this is giving complex solutions... but they're not reciprocals.. hmm

ganeshie8 (ganeshie8):

product of roots = -k/2 @idkihavenoidea

OpenStudy (whpalmer4):

well, of course a complex conjugate root pair can be reciprocal — just make \(a =0\) :-)

OpenStudy (anonymous):

@ganeshie8 when k is 18 the roots are -4.5, -1+i and -1-i => they are reciprocals!

OpenStudy (anonymous):

Oh - for cubics the product is \(-d/a\). Then it's \(-k/2\), yes.

OpenStudy (whpalmer4):

@BarbaraKara sorry, those are not reciprocals...

ganeshie8 (ganeshie8):

yeahh lol xD, only when a = 0 i guess... but why k=18 is not giving reciprrocal roots...

OpenStudy (anonymous):

whoops - conjugates i mean..

OpenStudy (anonymous):

\[\dfrac{1}{1+i} = \dfrac{1-i}{(1 + i)(1 - i)} = \dfrac{1 - i}{2}\]So \(1+i\) and \(0.5 - 0.5 i\) are reciprocals of each other. Just an example.

OpenStudy (whpalmer4):

yes, but they aren't a conjugate pair...

OpenStudy (anonymous):

We may, thus, plug in \(-k/2\) into our equation and try to solve for \(k\).

OpenStudy (whpalmer4):

and any polynomial with only real coefficients will only have complex conjugate roots

OpenStudy (anonymous):

Yes, I'm aware of that... just dropping an idea.

OpenStudy (whpalmer4):

okay, thought you were responding to this: "whether complex root pair can be reciprocals of each other"

ganeshie8 (ganeshie8):

looks like ive entered the equaitons incorrectly into wolfram, 1 sec -.-

OpenStudy (whpalmer4):

in any case, yes, plugging \(x = -k/2\) into \(2x^3+13x^2+22x + k = 0\) does give you an equation with 3 roots, one of which is the one that solves the problem.

OpenStudy (anonymous):

lol

OpenStudy (whpalmer4):

\[-\frac{k^3}{4}+\frac{13 k^2}{4}-10 k=0\]

ganeshie8 (ganeshie8):

wow ! yes, pluggin x = -k/2 in the polynomial is the best thing to do here xD

OpenStudy (anonymous):

The sum of roots is \(-13/2=-6.5\). If we're getting \(k = 5\) or \(k =8\) then the possible roots are \(-2.5\) and \(-4\).

OpenStudy (anonymous):

Ok but as @whpalmer4 says, "the "wrong" value corresponds to a value of k=5" right?

OpenStudy (anonymous):

Try to plug both \(-2.5\) and \(-4\) in.

ganeshie8 (ganeshie8):

both k = 5 and k = 8 looks like working

ganeshie8 (ganeshie8):

plug k = 5 and see if u get some a reciprocal pair or not..

OpenStudy (anonymous):

hmmm

OpenStudy (whpalmer4):

No, roots of \[2x^3+13x^2+22x+5=0\] are \(x = -\frac{5}{2}, x = -2\pm\sqrt{3}\)

OpenStudy (whpalmer4):

oops, sorry, @BarbaraKara don't read that, only @ganeshie8 :-)

OpenStudy (anonymous):

ok no worries :)

OpenStudy (anonymous):

Ahh, irrational conjugates. I had examples of them.

ganeshie8 (ganeshie8):

\(\large -2+\sqrt{3} = \dfrac{1}{-2-\sqrt{3}}\) @whpalmer4

OpenStudy (anonymous):

Oooh, thats interesting!

OpenStudy (whpalmer4):

I take back my earlier comment about 5 not working!

OpenStudy (anonymous):

thats all good :)

ganeshie8 (ganeshie8):

trickery tricky this is lol k= 5 gives : \(\large \dfrac{-5}{2}, ~-2 - \sqrt{3}, ~\dfrac{1}{-2-\sqrt{3}}\)

OpenStudy (whpalmer4):

However, \(k=8\) is clearly the superior solution, as both roots and \(k\) are multiples of each other, as previously mentioned :-)

ganeshie8 (ganeshie8):

hahah thats called solution biasing :P

OpenStudy (whpalmer4):

Though \(k=5\) is better for the head-scratching metric :-)

ganeshie8 (ganeshie8):

True that ! xD

OpenStudy (whpalmer4):

One wonders if the original author of the problem realized that there are two valid solutions...

OpenStudy (anonymous):

....solution bias, thats a first! haha

OpenStudy (anonymous):

knowing the trickster that is my teacher, yes she would know...

OpenStudy (whpalmer4):

perhaps it would be safer to say "realized that the solution is not unique" :-)

OpenStudy (whpalmer4):

then you should definitely submit the k=5 answer and see if you get called on it!

OpenStudy (anonymous):

will do!

ganeshie8 (ganeshie8):

you're wlcme \(\huge \color{red}{\ddot\smile}\)

OpenStudy (anonymous):

\( \color{red} {THANKS} \color{orange} {~HEAPS~} \color{green} {EVERYONE:~} \color{blue} {You~all~make~the~world~a~better~place!} \)

OpenStudy (anonymous):

Except for me, I just waste tonnes of oxygen everyday :p

OpenStudy (anonymous):

haha! - okay then..

OpenStudy (anonymous):

\( \huge \color{red}{\ddot\smile}~~ nice ~one! \) @ganeshie8

OpenStudy (anonymous):

...i'll stop now, I promise! haha - farewell, until my next question...

ganeshie8 (ganeshie8):

wow ! u became good at copy-pasting others latex code so quick ^_^

OpenStudy (anonymous):

well i was taught by the best wasn't I? haha

OpenStudy (anonymous):

Do I get to do the next question?

OpenStudy (anonymous):

why so keen @idkihavenoidea ?

ganeshie8 (ganeshie8):

Oh my xD ! that motivates me to try ur other questions lol :)

OpenStudy (anonymous):

I've been studying a similar topic

OpenStudy (anonymous):

Ooh awesome! I have lots more! hmmm let me see.....

ganeshie8 (ganeshie8):

okay, gtg.. . have a good day !

OpenStudy (anonymous):

Don't leave!

OpenStudy (anonymous):

Okay, here is one I had trouble with: but first, are u familiar with quotients and remainders @idkihavenoidea ?

OpenStudy (whpalmer4):

it's 3AM here, and I need to be up early, so you'll have to make do without me. I'll take a look in the morning and see if anyone has left me anything tractable to solve :-)